OpenStudy (anonymous):

If f(5)=3 and f'(5)=-2, find the derivative of x^2 f(x) at x=5.

OpenStudy (anonymous):

(x^2f(x))'=2xf(x)+x^2 f'(x) x=5 so = 2.5.f(5)+5^2 . f'(5)= 30-50=-20